A lactic acid solution is prepared by dissolving 0.67 g of lactic acid (HC3H5O3) in 100.00 mL of water. The pH is found to be 2.50. Calculate the Ka for lactic acid.
Answer #1

Molar mass of C3H6O3,

MM = 3*MM(C) + 6*MM(H) + 3*MM(O)

= 3*12.01 + 6*1.008 + 3*16.0

= 90.078 g/mol

mass(C3H6O3)= 0.67 g

use:

number of mol of C3H6O3,

n = mass of C3H6O3/molar mass of C3H6O3

=(0.67 g)/(90.08 g/mol)

= 7.438*10^-3 mol

volume , V = 1*10^2 mL

= 0.1 L

use:

Molarity,

M = number of mol / volume in L

= 7.438*10^-3/0.1

= 7.438*10^-2 M

use:

pH = -log [H+]

2.5 = -log [H+]

[H+] = 3.162*10^-3 M

HC3H5O3 dissociates as:

HC3H5O3 -----> H+ + C3H5O3-

7.438*10^-2 0 0

7.438*10^-2-x x x

Ka = [H+][C3H5O3-]/[HC3H5O3]

Ka = x*x/(c-x)

Ka = 3.162*10^-3*3.162*10^-3/(0.07438-3.162*10^-3)

Ka = 1.404*10^-4

Answer: 1.40*10^-4